In this question, we are going to derive the Surveyor’s Formula for finding the area of a polygon in terms of its coordinates.(a) If C is the line segment from (x1, y1) to (x2, y2) and ω=−ydx+xdy is a differential 1-form, show that ∫Cω=x1y2−x2y1 (b) If the vertices of a polygon (in counterclockwise order) are (x1, y1),(x2, y2), . . .(xn, yn),show that the area of the polygon is A=12[(x1y2−x2y1) + (x2y3−x3y2) +. . .(xn−1yn−xnyn−1) + (xny1+x1yn)] (c) Find the area of a pentagon with vertices (0,0),(2,1),(1,3),(0,2),(−1,1).
1
Expert's answer
2019-11-18T11:39:42-0500

a.Parametrize the straight line C by"x = x_1 + t(x_2 \u2212 x_1)" ,

"y = y_1 + t(y_2 \u2212 y_1),"

where t goes from "t = 0" to "t = 1" . Therefore

"\oint ydx+xdy=\int_{0} ^{1} −(y_1 + t(y_2 − y_1))(x_2 − x_1)dt +\int_{0} ^{1} (x_1 + t(x_2 − x_1))(y_2 − y_1)dt"

"= \u2212y_1(x_2 \u2212 x_1) \u2212(y_2 \u2212 y_1)(x_2 \u2212 x_1)\/2+ x1(y_2 \u2212 y_1) + (x_2 \u2212 x_1)(y_2 \u2212 y_1)"

"= \u2212y_1(x_2 \u2212 x_1) + x_1(y_2 \u2212 y_1) = x_1y_2 \u2212 x_2y_1" (Proved)

b.Consider two consecutive vertices"(x_i, y_i)" and"(x_{i+1}, y_{i+1})" . Let Ci be the line segment joining "(x_i,y_i)" to "(x_{i+1}, y_{i+1})" ", i = 1, . . . , n," with the convention that"(x_{n+1}, y_{n+1}) = (x_1, y_1)" . Let us compute

"\oint (xdy-ydx)"

To do so, we can parametrize "C_i" as

"x_i=(1-t)x_i + tx_{i+1}"

"y_i= (1 \u2212 t)y_i + ty_{i+1},"

"dx = (x_{i+1} \u2212 x_i) dt"

"dy = (y_{i+1 }\u2212 y_i) dt"

"\oint (xdy-ydx)= \int_{0} ^{1} ((1-t)x_i + tx_{i+1}).(y_{i+1 }− y_i) dt- \int_{0} ^{1} ((1 − t)y_i + ty_{i+1}). (x_{i+1} − x_i) dt"

"=\int_{0} ^{1}x_iy_{i+1}-x_i+2y_i"

"=x_iy_{i+1}-x_{i+1}" -----(1)

Denoting by D the region enclosed by the polygon

We know "A=1/2\smallint _{dD} ( xdy-ydx)"

"=1/2\Sigma _{i=1}^n (\smallint _{C_i}(xdy-ydx))"

Using (1): we get the desired result

"A=1\/2[(x_1y_2\u2212x_2y_1) + (x_2y_3\u2212x_3y_2) +. . .(x_{n\u22121}y_n\u2212x_ny_{n\u22121}) + (x_ny_1+x_1y_n)]"

(Proved)

(c)"Area=1\/2[(0-0)+(2.3-1.1)+(1.2-0)+(0-(-2))+(0-0)]"

"=1\/2(5+2+2)=9\/2" (Answer) using formula proved in (b)